OpenStudy (anonymous):

What is the simplified form of 18 times g cubed times h to the ninth power all over 6 times g to the sixth power times h to the fifth power.

OpenStudy (anonymous):

\[\large \frac{18g^3 h^9}{6g^6h^{15}}\] likek that?

OpenStudy (anonymous):

yeah but its a 5 next to the h not a 15

OpenStudy (anonymous):

\[\large \frac{18g^3 h^9}{6g^6h^{5}}\]

OpenStudy (anonymous):

18 divide by 6 is 3 so you will have a 3 up top

OpenStudy (anonymous):

\[\large \frac{18g^3 h^9}{6g^6h^{5}}\large =\frac{6h^{9-5}}{g^{6-3}}\] subtract the smaller from the larger if your job is to rewrite using only positive exponents

OpenStudy (anonymous):

damn that was wrong, sorry numerator should have a 3 not a 6

OpenStudy (anonymous):

\[\large \frac{18g^3 h^9}{6g^6h^{5}}\large =\frac{3h^{9-5}}{g^{6-3}}=\frac{3h^4}{g^3}\]

OpenStudy (anonymous):

so basically you just divide everything? and subtract the exponent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not very hard, is it? really what you are doing is cancelling the common factors

OpenStudy (anonymous):

thank you so much! this will help me with my other problems (: